Calculus II, Spring 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 10

Assigned April 30, 2009.

1. Read page 884, 900-901 on how to get limits of integration. This is the most important topic in this Chapter.

2. Section 14.1: Problems 11, 19, 23, 27, 31, 33, 37, 39, 45.

3. Section 14.3: Problems 9, 13, 15, 23.

4. Section 14.4: Problems 3, 11, 13, 29, 27, 35.